OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20

OpenStudy (lalaly):

g(x)=20 --> 2(x-4)=20 now solve for x

OpenStudy (anonymous):

20=2(x-4) 10=x-4 14=x